# Infinite quantum twisting at the Cauchy horizon of rotating black holes

### v1

@@ I Introduction
The validity of sCC has been investigated in many works, starting with Poisson and Israel (1989, 1990); Ori (1991) who accumulated suggestive evidence in favor of the hypothesis. One question concerns the exact nature of the singularity. Recent mathematical works on the full Einstein equations Dafermos and Luk (2017) show that tidal distortions at the singularity remain finite for sufficiently small initial perturbations  [^1] For essentially nonlinear effects in spherical symmetry, see Li and Van de Moortel (2023)., but one expects that tidal forces suitably diverge in such a way as to make the metric inextendible as a weak solution to the Einstein equations. According to a formulation by Christodoulou (2009), this means that first derivatives of the metric ought to fail to be locally square integrable at the IH. In terms of scalar test-fields, it means that some component $T_{\mu\nu}$ of the matter stress-energy should fail to be locally integrable there. Although a proof of this strong form of sCC is still missing in the context of the full Einstein-matter equations, there is a large body of evidence in the context of the test field approximation of the Einstein equations, and via numerical approaches Mellor and Moss (1990, 1992); Brady et al. (1998); Dafermos and Shlapentokh-Rothman (2017); Luk and Oh (2017); Cardoso et al. (2018a); Dias et al. (2019, 2018); Costa et al. (2018); Luna et al. (2019).
@@
The rest of this Letter is organized as follows. Secs. II and III respectively introduce KdS spacetime and scalar field quantum field theory on KdS. The formula for the RSET at the IH is derived in Sec. IV. Sec. V contains our numerical results for the RSET. We summarize the results in Sec. VI. Throughout, we set $\hbar=c=G_{N}=1$.
@@ II Geometric setup
Here, $\Delta$ is a 3-dimensional surface tangent to $\xi$ bounding ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$, and $\,\text{d}\Sigma^{\alpha}$ is an appropriately oriented covariant integration element on $\Delta$. We are interested in the case where ${\mathcal{S}}_{2}$ approaches the future IH, ${\mathcal{H}}^{R}_{-}$, at constant value of $u$, while ${\mathcal{S}}_{1}$, located at the same constant $u$, stays at some constant value $V_{-}<0$. We will assume that, to estimate quantum effects, the stress tensor can be replaced in the small backreaction regime by the RSET ${\left\langle\hat{T}_{\alpha\beta}\right\rangle}_{\mathrm{U}}$; we denote its value at the IH by ${\left\langle\hat{T}_{\alpha\beta}(\theta)\right\rangle}^{\rm IH}_{{\mathrm{U}}}$.
@@ III The scalar quantum field
However, the Boulware modes $\phi^{\mathrm{in}/\mathrm{up},{\mathrm{I}}}_{\omega\ell m}$ (with reflection and transmission coefficients $\mathcal{R}^{\mathrm{in}/\mathrm{up},{\mathrm{I}}}_{\omega\ell m}$ and $\mathcal{T}^{\mathrm{in}/\mathrm{up},{\mathrm{I}}}_{\omega\ell m}$) and $\phi^{\mathrm{in}/\mathrm{up},{\mathrm{II}}}_{\omega\ell m}$, defined in Sup , are easier to calculate than the Unruh modes. By expressing the Unruh modes in terms of Boulware modes similarly as in Klein and Zahn (2021); Zilberman et al. (2022b), we obtain for the anti-commutator in ${\mathrm{II}}$:
@@ IV The stress-energy tensor
where $D_{\alpha\beta^{\prime}}$ is a bi-differential operator given in Sup and $x$ is any point where both states are defined.
We construct the comparison state similarly to Hollands et al. (2020a) for RN: we modify the metric in $r<\delta<r_{-}$, with $0<\delta\ll r_{-}$, replacing the singularity at $r=0$ Combined with Eqs. (9) and (10), this yields Sup :
@@ V Numerical results
We plot in Fig. 3 the angular momentum current at the IH as a function of $\cos\theta$ for $\Lambda=1/30$ and $a=0.95$, $0.975$, $1$, and $a_{\text{max}}-1/200$, where $a_{\text{max}}\approx 1.012$. In Sup we also plot the angular momentum current at $\cos\theta=0.9$ as a function of $a$ for different values of $\Lambda$. Our first main result is that ${\left\langle T_{v\varphi_{-}}\right\rangle}_{\mathrm{U}}^{\rm IH}$ can have different signs depending on the latitude $\theta$.
@@
Figure 4: ${\left\langle\hat{T}_{vv}\right\rangle}_{{\mathrm{U}}}^{\rm IH}$ for $\theta=0$, $\Lambda=\frac{1}{270}$ as a function of $a$. The vertical line indicates $a_{\text{max}}$. The inset zooms in near extremality.
@@ VI Conclusions
As we argue in Sup , if we extrapolate the qualitative features of our findings to a setting with backreaction, then topological spheres approaching the IH undergo infinite expansion in some parts, and contraction in others. On a latitude separating such regions, we have a diverging relative twisting of neighboring latitudes and/or a twisting of perpendicular incident lightrays towards or against the $\varphi_{-}$-direction, Fig. 1.

### v2

@@ I Introduction
The validity of sCC has been investigated in many works, starting with Poisson and Israel (1989, 1990); Ori (1991) who accumulated suggestive evidence in favor of the hypothesis. One question concerns the exact nature of the singularity. Recent mathematical works on the full Einstein equations Dafermos and Luk (2017) show that tidal distortions at the singularity remain finite for sufficiently small initial perturbations  [^1] For essentially nonlinear effects in spherical symmetry, see Li and Van de Moortel (2023)., but one expects that tidal forces suitably diverge in such a way as to make the metric inextensible as a weak solution to the Einstein equations. According to a formulation by Christodoulou (2009), this means that first derivatives of the metric ought to fail to be locally square integrable at the IH. In terms of scalar test-fields, it means that some component $T_{\mu\nu}$ of the matter stress-energy should fail to be locally integrable there. Although a proof of this strong form of sCC is still missing in the context of the full Einstein-matter equations, there is a large body of evidence in the context of the test field approximation of the Einstein equations, and via numerical approaches Mellor and Moss (1990, 1992); Brady et al. (1998); Dafermos and Shlapentokh-Rothman (2017); Luk and Oh (2017); Cardoso et al. (2018a); Dias et al. (2019, 2018); Costa et al. (2018); Luna et al. (2019).
@@
The rest of this Letter is organized as follows. Secs. II and III respectively introduce KdS spacetime and scalar field quantum field theory on KdS. The formula for the RSET at the IH is derived in Sec. IV. Sec. V contains our numerical results for the RSET. We summarize the results in Sec. VI. App. A provides details on the background spacetime, the calculations of the RSET and the twisting of spheres. Throughout, we set $\hbar=c=G_{N}=1$.
@@ II Geometric setup
Here, $\Delta$ is a 3-dimensional surface tangent to $\psi^{\alpha}$ bounding ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$, and $\,\text{d}\Sigma^{\alpha}$ is an appropriately oriented covariant integration element on $\Delta$. We are interested in the case where ${\mathcal{S}}_{2}$ approaches the future IH, ${\mathcal{H}}^{R}_{-}$, at constant value of $u$, while ${\mathcal{S}}_{1}$, located at the same constant $u$, stays at some constant value $V_{-}<0$. We will assume that, to estimate quantum effects, the stress tensor can be replaced in the small backreaction regime by the RSET ${\left\langle\hat{T}_{\alpha\beta}\right\rangle}_{\mathrm{U}}$; we denote its value at the IH by ${\left\langle\hat{T}_{\alpha\beta}(\theta)\right\rangle}^{\rm IH}_{{\mathrm{U}}}$.
@@ III The scalar quantum field
However, the Boulware modes $\phi^{\mathrm{in}/\mathrm{up},{\mathrm{I}}}_{\omega\ell m}$ (with reflection and transmission coefficients $\mathcal{R}^{\mathrm{in}/\mathrm{up},{\mathrm{I}}}_{\omega\ell m}$ and $\mathcal{T}^{\mathrm{in}/\mathrm{up},{\mathrm{I}}}_{\omega\ell m}$) and $\phi^{\mathrm{in}/\mathrm{up},{\mathrm{II}}}_{\omega\ell m}$, defined in App. A, are easier to calculate than the Unruh modes. By expressing the Unruh modes in terms of Boulware modes similarly as in Klein and Zahn (2021); Zilberman et al. (2022b), we obtain for the anti-commutator in ${\mathrm{II}}$:
@@ IV The stress-energy tensor
where $D_{\alpha\beta^{\prime}}$ is a bi-differential operator given in App. A and $x$ is any point where both states are defined.
We construct the comparison state similarly to Hollands et al. (2020a) for RN: we modify the metric in $r<\delta<r_{-}$, with $0<\delta\ll r_{-}$, replacing the singularity at $r=0$. Combined with Eqs. (9) and (10), this yields (App. A):
@@ V Numerical results
We plot in Fig. 3 the angular momentum current at the IH as a function of $\cos\theta$ for $\Lambda=1/30$ and $a=0.95$, $0.975$, $1$, and $a_{\text{max}}-1/200$, where $a_{\text{max}}\approx 1.012$. In App. A, we also plot the angular momentum current at $\cos\theta=0.9$ as a function of $a$ for different values of $\Lambda$. Our first main result is that ${\left\langle T_{v\varphi_{-}}\right\rangle}_{\mathrm{U}}^{\rm IH}$ can have different signs depending on the latitude $\theta$.
@@
Figure 4: ${\left\langle\hat{T}_{vv}\right\rangle}_{{\mathrm{U}}}^{\rm IH}$ for $\theta=0$, $\Lambda=\frac{1}{270}$ as a function of $a$. The vertical line indicates $a_{\text{max}}$. The inset is a closeup near extremality.
@@ VI Conclusions
As we argue in App. A, if we extrapolate the qualitative features of our findings to a setting with backreaction, then topological spheres approaching the IH undergo infinite expansion in some parts, and contraction in others. On a latitude separating such regions, we have a diverging relative twisting of neighboring latitudes and/or a twisting of perpendicular incident lightrays towards or against the $\varphi_{-}$-direction, Fig. 1.